OpenStudy (anonymous):

find the equation ofthe locus of a point which moves in the plane so that its distance from (-2,0) is always three times its distance fromt he origin

OpenStudy (amistre64):

thats a parabola

OpenStudy (anonymous):

the question after it says shows that the locus is a circle...

OpenStudy (amistre64):

the distance from its single foci is 3times its distance to the origin .... might be eliptic

OpenStudy (amistre64):

i cant tell :)

OpenStudy (saifoo.khan):

hey amistre.

OpenStudy (amistre64):

yes? :)

OpenStudy (saifoo.khan):

if v hav a graph of x = 1, then wht will b y ?

OpenStudy (amistre64):

y will be lonely

OpenStudy (saifoo.khan):

but i cant write this LONELY thing,

OpenStudy (amistre64):

\(\color{#f2e526}{\text{i}}\)\(\color{#3f5611}{\text{f}}\)\(\color{#b55857}{\text{ }}\)\(\color{#52da8d}{\text{x}}\)\(\color{#43eb8c}{\text{=}}\)\(\color{#6ed573}{\text{1}}\)\(\color{#2fcdae}{\text{;}}\)\(\color{#2dadc9}{\text{ }}\)\(\color{#a2bf6f}{\text{y}}\)\(\color{#3bddb5}{\text{ }}\)\(\color{#fb28fd}{\text{w}}\)\(\color{#35b173}{\text{i}}\)\(\color{#da931a}{\text{l}}\)\(\color{#e1c579}{\text{l}}\)\(\color{#d2396b}{\text{ }}\)\(\color{#5398c5}{\text{a}}\)\(\color{#eefeac}{\text{l}}\)\(\color{#3adb86}{\text{w}}\)\(\color{#aabcb4}{\text{a}}\)\(\color{#d9ac7d}{\text{y}}\)\(\color{#bd46ed}{\text{s}}\)\(\color{#fc8c8d}{\text{ }}\)\(\color{#1c55a2}{\text{b}}\)\(\color{#ffa143}{\text{e}}\)\(\color{#1de66e}{\text{ }}\)\(\color{#bbf359}{\text{l}}\)\(\color{#d5f6ff}{\text{o}}\)\(\color{#cedad1}{\text{n}}\)\(\color{#4c94be}{\text{e}}\)\(\color{#f7ad29}{\text{l}}\)\(\color{#49a7f2}{\text{y}}\)\(\color{#864fb7}{\text{}}\)

OpenStudy (anonymous):

lol okay...

OpenStudy (amistre64):

y not exist on a vertical line do it?

OpenStudy (saifoo.khan):

no.

OpenStudy (anonymous):

umm so...circle...or no...where'd amistre go?

OpenStudy (saifoo.khan):

but wht should i write?

OpenStudy (amistre64):

location of a point which moves in the plane its distance from (-2,0) is always 3x distance from (0,0)

OpenStudy (anonymous):

yes...

OpenStudy (amistre64):

what is its distance if it IS the origin

OpenStudy (anonymous):

[\sqrt{(x-0)^2+(y-0)^2}]

OpenStudy (anonymous):

for my final answer i got 2x^2-4x-4

OpenStudy (amistre64):

the point (-3,0) is on this path; and so is (-1.5 , 0 )

OpenStudy (amistre64):

right?

OpenStudy (anonymous):

right!

OpenStudy (anonymous):

and we did that because thats our new second point?

OpenStudy (amistre64):

thats just one point that is 3 times the distance to the origin as it is to the -2,0

OpenStudy (amistre64):

i just tried to figure it for a point above and below the point; but my angle comes back as 0 or 180

OpenStudy (anonymous):

this question is very interesting...

OpenStudy (amistre64):

its cause I tried to force it to be 3 and 1 lol instead of a ration

OpenStudy (amistre64):

if i did it right; we get points above and below it at a distance of 1/sqrt(2)

OpenStudy (amistre64):

well the good news is it aint a circle

OpenStudy (anonymous):

thats bad news lol...

OpenStudy (amistre64):

lets flip it over to the positive side from 0 to 2 is our origin and point; 3,0 is 1 away from 2 and 3 away fromm 0; its good 1.5 is 3/2 from 0 and 1/2 from 2 so we have 2 points (1.5,0) and (3,0) that satisfy our misinterpretations

OpenStudy (anonymous):

u are over complicating it

OpenStudy (amistre64):

the center of these 2 points is of course 2.25; if this was a circle then the point that is above and below it would match our problem

OpenStudy (amistre64):

overcomplicating is what I do when im using what i know to get to where i dont know

OpenStudy (anonymous):

\[\sqrt{(x+2)^2 +y^2} = 3 \sqrt{ x^2+y^2}\]

OpenStudy (anonymous):

square both sides expand, group and complete the squares

OpenStudy (amistre64):

.75 would be our radius by my conjecture; so the point (2.25, .75) would have to pythag to = 9

OpenStudy (anonymous):

(x^2 +4x +4 +y^2 = 9 ( x^2 +y^2)

OpenStudy (amistre64):

well, not 9; but 3(.75 ) to be good

OpenStudy (anonymous):

8x^2 +8y^2 -4x -4 =0

OpenStudy (anonymous):

divide through by 4

OpenStudy (anonymous):

no 8

OpenStudy (anonymous):

x^2 -(1/2)x +y^2 = (1/2)

OpenStudy (anonymous):

(x-(1/4) )^2 +y^2 = (1/2) + (1/16)

OpenStudy (anonymous):

(x-(1/4))^2 +y^2 = 9/16

OpenStudy (anonymous):

i got the same thing but i forgot tot square my 3

OpenStudy (anonymous):

circle ecentre ( (1/4) , 0 ) radius 3/4

OpenStudy (amistre64):

if hes right; id go with it cause I cant get nothing beating my head against the wall lol

OpenStudy (anonymous):

lol aww thanks so much amistre...

OpenStudy (anonymous):

so elece you used method of completing of the squares to find the centre and radius

OpenStudy (anonymous):

THANKS SO MUCH ELECENGINEER!!! and AMISTRE!!!

OpenStudy (anonymous):

9/16 is not 3/4

OpenStudy (anonymous):

circle is x^2 +y^2 = r^2

OpenStudy (anonymous):

sqrt(9/16)= 3/4